OpenStudy (anonymous):

in 2006 the number of tickets sold was 210. the ratio; (no. of adult tickets: no. of student tickets) was 23:19. how many tickets were sold?

OpenStudy (lgbasallote):

use 23x + 19x = 210 can you solve for x?

OpenStudy (anonymous):

wat?

OpenStudy (anonymous):

sorry cz i forgot to medal you in the previous question :D

OpenStudy (lgbasallote):

the ratio is 23:19 then it equals to 210 therefore 23x + 19x = 210 x is the number of tickets

OpenStudy (lgbasallote):

solve for x

OpenStudy (anonymous):

how??

OpenStudy (lgbasallote):

what is 23 + 19?

OpenStudy (anonymous):

42

OpenStudy (lgbasallote):

so 23x + 19x is?

OpenStudy (anonymous):

42

OpenStudy (lgbasallote):

42x

OpenStudy (lgbasallote):

so 23x + 19x = 210 42x = 210 now divide both sides by 42 btw...is the question right??? how many tickets were sold???